OpenStudy (anonymous):

Use substitution to solve the system of equations. {3x = 6y + 12 {18y = 9x -32 A. (2, -1) B. (38/9, 1/9) C. (6, 1) D. No solution (I think it's this one)

OpenStudy (anonymous):

(x,y) plug the number that corresponds to x and y into the equation. For example, answer A has (2,-1), so plug in 2 for x and -1 for y into your equation.

OpenStudy (anonymous):

Does that make sense to you?

OpenStudy (anonymous):

So, would it be A because: 3x = 6y + 12 3(2) = 6(-1) + 12 6 = 6

OpenStudy (anonymous):

Yes, it does work for the first equation, but it should also work for the second equation for it to be right.

OpenStudy (anonymous):

If I plug it into the second equation I get: 18y = 9x - 32 18(-1) = 9(2) - 32 -18 = 18 - 32 -18 = -14 so A is not right.

OpenStudy (anonymous):

you got it. You have to do the same for the other values until you find a (x,y) that works in both. If you cannot find one that works for both, then the answer would be no solution.

OpenStudy (anonymous):

I did C and it doesn't work either, but I'm not sure how to do B because it's fractions. Can you help?

OpenStudy (anonymous):

I would use a calculator since the point of this problem is to learn to do system of equations, not to learn fractions.

OpenStudy (anonymous):

It sounds like you understand system of equations since you worked out A and C to be false. So, your initial guess of no solution was right.